OpenStudy (anonymous):

Need to find the perpidicular line of 7x +5y = -3

OpenStudy (anonymous):

The slope of a line perpendicular to a line of slope m is -1/m. What's the slope of the given line?

OpenStudy (anonymous):

First simplify your formula to standard y=mx+b format: 7x+5y=-3 5y=-7x-3 y=(-7/5)x-(3/5) Now the perpendicular slope is -1/m so: -1/(-7/5)=5/7 So your answer is y=(5/7)x If you want it to have the same intercept as your original formula then it would be y=(5/7)x-3